OpenStudy (anonymous):

Find a value for a so that the function f(x)={4-x^2, x<-1 and ax^2-1, x>=-1 is continuous.

OpenStudy (dumbcow):

evaluate at x = -1.....4-(-1)^2 = 4-1 = 3..............a(-1)^2-1 = a-1..................the 2 parts should be equal at x=-1 for f(x) to be continuous.............a-1 = 3 --> a = 4